# Mirror Speculative Decoding: Breaking the Serial Barrier in LLM Inference

### v1

@@ 1 Introduction
Autoregressive (AR) large language models (LLMs) have achieved state-of-the-art performance across a wide spectrum of natural language processing (NLP) tasks, yet their decoding latency remains a fundamental bottleneck, particularly for real-time applications such as interactive dialogue, code generation, and on-device assistants  (Brown et al., 2020; Pope et al., 2023). Speculative decoding (SD) has emerged as a promising paradigm to mitigate this limitation by coupling a lightweight draft model with a larger, high-fidelity target model (Leviathan et al., 2023; Chen et al., 2023). In the canonical two-model SD framework, the draft model generates candidate tokens which are then verified by the target model in a serial pipeline. While this approach reduces the number of target model invocations, the sequential dependency between draft and target stages limits achievable speedups. Recent works attempt to relax the serial constraints by equipping the target itself with speculative capacity. Medusa  (Cai et al., 2023) equips the target with parallel decoding heads, while EAGLE  (Li et al., 2024a) introduces a dedicated speculation layer. However, the same trade-off remains: larger speculative modules improve acceptance at the cost of higher draft construction latency, while smaller ones reduce overhead but lower acceptance and limit speedup. A detailed discussion of related approaches is provided in Appendix A.
@@
Modern System on Chip (SoC) architectures increasingly feature heterogeneous compute units that combine general-purpose CPUs with specialized accelerators such as GPUs and dedicated neural processing units (NPUs) (Jouppi et al., 2021; Intel Corporation, 2023; Advanced Micro Devices (2023), AMD). This design trend enables efficient partitioning of workloads across compute substrates optimized for different performance and power trade-offs. For instance, Apple’s M-series chips integrate a high-throughput GPU and a dedicated Apple Neural Engine (ANE)  (Apple Inc., 2023a; b). Similarly, Qualcomm’s Snapdragon 8 Gen 3 features an Adreno GPU alongside a Hexagon NPU optimized for mixed-precision inference (Qualcomm Technologies Inc., 2023). This architectural heterogeneity motivates a division-of-labor strategy for speculative decoding, wherein the draft model operates on the NPU exploiting its efficiency for approximate inference, while the target model executes on the GPU, which is better suited for high-fidelity, throughput-critical computation. Such partitioning leverages available NPU capacity and reduces contention on the GPU, thereby improving end-to-end latency in multi-accelerator deployments.
@@ 2 Speculative Decoding: Formalization and Limits
where $h_{t}\in\mathbb{R}^{H}$ is the next-token representation at position $m+t$, and $W\in\mathbb{R}^{|\mathcal{V}|\times H}$ is the output head mapping hidden states to vocabulary logits (Radford & Narasimhan, 2018; Vaswani et al., 2017). Scaling inference of such models often requires distributing computation across multiple devices via tensor parallelism, which partitions per-layer parameters across devices and aggregates partial results with collectives such as AllReduce (Hansen-Palmus et al., 2024; Li et al., 2024e). The per-token latency is then set by the critical path combining local compute and synchronizations.
Speculative decoding (SD). Speculative decoding augments a target model $f_{target}(\cdot\mid\cdot)$ with a computationally cheaper draft model $f_{draft}(\cdot\mid\cdot)$  (Leviathan et al., 2023; Chen et al., 2023). At step $t$, conditioned on the verified prefix $(x,y_{<t})$, the draft proposes a $\gamma$-token window
@@
Increasing draft capacity (larger $\gamma$, deeper/wider $f_{d}$) typically increases $\rho$ but also increases $T_{\text{draft}}$, while tiny drafts reduce $T_{\text{draft}}$ but suffer low $\rho$ (Leviathan et al., 2023; Chen et al., 2023).  Equation 5 exposes the core limitation: improvements in acceptance must compensate for the added draft latency, intrinsically coupling acceptance with latency.
@@ 3 Mirror Speculative Decoding
which is typically strongly correlated with the final distribution $p^{(N)}(\cdot\mid y_{<t},x)$  (Pal et al., 2023a). We designate an early-exit layer $\ell_{e}\in\{1,\dots,N-1\}$ and expose a low-bandwidth token channel:
@@ 3.2 Draft execution with Speculative Streaming
For the draft model $f_{d}$, we employ Speculative Streaming (SS) (Bhendawade et al., 2024), a speculative mechanism that verifies previously proposed tokens while generating new speculative tokens in the same forward pass using multi-stream attention. Applying SS to the target would modify its decoding dynamics and alter the final distribution $p^{(N)}(\cdot\mid y_{<t},x)$ (Bhendawade et al., 2024), breaking the lossless guarantee established in  Appendix B. In contrast, using SS on the draft accelerates speculation generation without changing acceptance semantics, since all commitments still require verification against the unchanged target. This design leverages SS precisely where it yields additional concurrency while preserving correctness (See  Appendix B).  Section D.2 illustrates the SS mechanism and compares draft-only speedups between vanilla and SS drafts.
@@
Within each Mirror-SD step, the draft must furnish a branch-complete speculative window of length $\gamma$ at the rendezvous ( Section 3.1). Under SS, a single draft internal step can emit $\eta_{j}\!\geq\!1$ tokens by verifying the prior proposal and predicting multiple future tokens in one pass (Bhendawade et al., 2024). Consequently, the number of draft steps $J$ required to materialize $\gamma$ tokens satisfies
@@ 3.3 Heterogeneous Sharding of Mirror-SD
The draft is trained with SPD architecture  (Kim et al., 2025). We divide the $N_{\mathrm{d}}$ layers into two contiguous segments. Within each segment we instantiate $G_{D}$ parallel tracks; track $g\in\{1,\dots,G_{D}\}$ is pinned to NPU $g$ and advances through all layers of its segment using a resident weight shard. There is no inter-NPU traffic inside a segment (See  Figure 2). At the segment boundary, all tracks perform a single global synchronization to re-align tensor partitions, and a second synchronization occurs at the end of the forward pass to assemble full-width logits for the main and lookahead streams. Each internal draft step executes two all-reduce collectives on activation shards while weights remain sharded. This replaces per-layer synchronization with a fixed two-collective cost, reducing latency and enabling more parameters to be sharded across NPUs. In practice, this expands draft capacity and improves acceptance rates $\rho(\gamma;\phi,\theta)$ without increasing critical-path latency.
@@ 4.1 Evaluation protocol
Datasets and tasks. We integrate our approach with the open-source SpecBench framework (Xia et al., 2024b) to ensure a fair, reproducible comparison against prior methods. SpecBench provides standardized prompts and pre/post-processing, sampling settings and released configs and seeds (Xia et al., 2024b). We report results on multi-turn interactive conversation (MT Bench), translation, summarization, mathematical reasoning, machine translation and retrieval-augmented generation (RAG). Context and generation lengths follow the SpecBench protocol (Xia et al., 2024b).
Models and baselines. We evaluate Mirror-SD on server-scale targets that are deployable in production inference stacks: Qwen3-14B and Qwen3-32B (Yang et al., 2025), Mistral-24B (Mistral, 2025), and OPT-66B (Zhang et al., 2022). For Qwen targets, we train a 0.6B-parameter draft with 2 segments and 8 tracks and deploy it on 8 NPUs as described in Section 3.3. For Mistral we train a 0.5B draft, and for OPT we train a 200M draft, both sharded as in Section 3.3 to optimize synchronization cost. All draft models are trained with SS objective described in (Bhendawade et al., 2024) on UltraChat  (Ding et al., 2023). Across all target models, drafts are launched from the mid-layer early exit ($\frac{1}{2}$ of total depth) with top-$\kappa$=8 under batch size 1. Please refer to Appendix E for the effects of early-exit depth and $\kappa$. Baselines include vanilla SD, Medusa (Cai et al., 2024), Hydra (Ankner et al., 2024b), EAGLE 2/3 (Li et al., 2024b; 2025a), Recycling (Luo et al., 2024), PLD (Saxena, 2023a), SpS (Joao Gante, 2023), REST (He et al., 2024), and Lookahead (Fu et al., 2023). All baselines have public implementations in SpecBench (Xia et al., 2024b), and we use the corresponding implementations.
@@
Serving configuration and reproducibility. Target models are distributed across eight M2 Ultra GPUs using Megatron-style tensor parallelism (Section 3.3), while the draft runs on eight NPUs (Apple Inc. (2023a)). All evaluations use a fixed batch size of 1 and speculative window length $\gamma{=}7$; please refer to Section D.1 for analysis of batching effects. The token channel transmits only the top-$\kappa$ token IDs and log-probabilities in bf16. For determinism, interconnects are pinned and frequency scaling is disabled. Timings include compute, collectives, and rendezvous overhead.
@@ 4.2 Tri-objective analysis with an MT-Bench diagnostic
Speculative decoding couples three quantities: the speculative window $\gamma$, the acceptance length $\mathbb{E}[A_{t}]=\gamma\,\rho(\gamma;\phi,\theta)$, and the per-speculation-step latency. In vanilla SD, enlarging $\gamma$ typically boosts acceptance but also increases draft construction time, yielding an upward-sloping latency curve. For Mirror-SD, the step latency follows the model in  Section 3.4 (Equation 10): as long as $T_{\text{draft}}^{\mathrm{gen}}(\gamma)\!\leq\!\Delta$ with $\Delta=T_{\text{target}}^{\,\ell_{e}+1:N}$, increasing $\gamma$ (and thus $\mathbb{E}[A_{t}]$) adds no marginal latency; once $T_{\text{draft}}^{\mathrm{gen}}(\gamma)\!>\!\Delta$, latency grows by the excess beyond $\Delta$. Acceptance semantics remain unchanged ( Appendix B). We validate these hypotheses on MT-Bench (Bai et al., 2024) by sweeping $\gamma$, measuring $\mathbb{E}[A_{t}]$ and the observed draft construction time, and comparing three methods that share the same target: (i) vanilla SD with autoregressive drafts from 12M to 1.7B parameters, (ii) Mirror-SD with a 0.60.6M draft, and (iii) Mirror-SD with a speculative-streaming draft (Section 3.2) of 0.60.6B.  Figure 3(a) places $\gamma$ on the x-axis, $\mathbb{E}[A_{t}]$ on the y-axis, and draft construction latency on the z-axis.
@@ Appendix A Related Works
The original speculative decoding paradigm accelerates autoregressive generation by pairing a small, fast draft model with a larger target model, which verifies proposed tokens (Chen et al., 2023; Leviathan et al., 2023). This approach achieves substantial wall-time savings whenever the draft is hardware-efficient and closely aligned with the target. Domain-specialized drafts trained via distillation further improve acceptance in task-specific settings (Hong et al., 2025). Recent variants explore parallelization strategies, such as batch-axis speculation (Sun et al., 2023b) and tree-structured drafts (Miao et al., 2023; Spector & Re, 2023), to raise acceptance rates and amortize draft cost.
@@
An alternative line of work removes the explicit draft model and equips the target itself with speculative capacity. Medusa predicts multiple tokens in parallel via extra heads (Cai et al., 2023), while Hydra enforces autoregressive coupling across those heads to raise acceptance (Ankner et al., 2024b). EAGLE introduces a dedicated speculation layer (Li et al., 2024a), with EAGLE-2 enabling dynamic tree retries (Li et al., 2024b) and EAGLE-3 moving to token-level prediction with multi-layer fusion (Li et al., 2025a). Prompt-lookup decoding (PLD) and Lookahead propose suffixes by retrieval rather than generation (Saxena, 2023a; Fu et al., 2023), which is effective when prefix–continuation correlations are strong. Recycling reduces wasted work by reusing intermediate activations when speculative branches are invalidated, instead of recomputing full forwards (Luo et al., 2024). Other recent advances include structured or retrieval-based decoding policies (Yi et al., 2024a; He et al., 2024). Across the single-model designs, speculative capacity is integrated into the target stack, so larger or wider modules increase acceptance but still add work on the target’s critical path; by contrast, Mirror-SD runs draft and target on heterogeneous devices and overlaps draft within the target’s suffix window, converting added draft capacity into acceptance gains without inflating per-step latency proportionally.
@@
Beyond speculation, a range of methods accelerate inference by adapting compute during decoding. CALM (Schuster et al., 2022) and related early-exit methods reduce cost by exiting tokens at shallow layers, while skip decoding (Corro et al., 2023) mitigates key-value cache mismatch via position-dependent layer skipping. Mixture-of-Depths (MoD) (Raposo et al., 2024) routes only a subset of tokens through full blocks, yielding non-uniform FLOP allocation. Other strategies include token merging (Bolya et al., 2023) to reduce sequence length dynamically, adaptive span models (Sukhbaatar et al., 2019) that learn context windows per token, and CoLT5 (Ainslie et al., 2023) which routes tokens through heavy or light pathways. More recently, M2R2 (Bhendawade et al., 2025) introduces accelerated residual streams to improve early alignment and efficiency. Together, these approaches trade fixed per-token compute for dynamic allocation, complementing speculative decoding’s strategy of parallelizing token generation.

### v2

@@ 1 Introduction
Autoregressive (AR) large language models (LLMs) have achieved state-of-the-art performance across a wide spectrum of natural language processing (NLP) tasks, yet their decoding latency remains a fundamental bottleneck, particularly for real-time applications such as interactive dialogue, code generation, and on-device assistants  (brown2020language; pope2023efficiently). Speculative decoding (SD) has emerged as a promising paradigm to mitigate this limitation by coupling a lightweight draft model with a larger, high-fidelity target model (leviathan2023fast; chen2023accelerating). In the canonical two-model SD framework, the draft model generates candidate tokens which are then verified by the target model in a serial pipeline. While this approach reduces the number of target model invocations, the sequential dependency between draft and target stages limits achievable speedups. Recent works attempt to relax the serial constraints by equipping the target itself with speculative capacity. Medusa  (medusa) equips the target with parallel decoding heads, while EAGLE  (eagle) introduces a dedicated speculation layer. However, the same trade-off remains: larger speculative modules improve acceptance at the cost of higher draft construction latency, while smaller ones reduce overhead but lower acceptance and limit speedup. A detailed discussion of related approaches is provided in Appendix A.
@@
Modern System on Chip (SoC) architectures increasingly feature heterogeneous compute units that combine general-purpose CPUs with specialized accelerators such as GPUs and dedicated neural processing units (NPUs) (jouppi2021ten; intel_ultra_2023; amd_ryzen_ai_2023). This design trend enables efficient partitioning of workloads across compute substrates optimized for different performance and power trade-offs. For instance, Apple’s M-series chips integrate a high-throughput GPU and a dedicated Apple Neural Engine (ANE)  (apple_m2_ultra; apple_m3_2023). Similarly, Qualcomm’s Snapdragon 8 Gen 3 features an Adreno GPU alongside a Hexagon NPU optimized for mixed-precision inference (qualcomm_snapdragon_2023). This architectural heterogeneity motivates a division-of-labor strategy for speculative decoding, wherein the draft model operates on the NPU exploiting its efficiency for approximate inference, while the target model executes on the GPU, which is better suited for high-fidelity, throughput-critical computation. Such partitioning leverages available NPU capacity and reduces contention on the GPU, thereby improving end-to-end latency in multi-accelerator deployments.
@@ 2 Speculative Decoding: Formalization and Limits
where $h_{t}\in\mathbb{R}^{H}$ is the next-token representation at position $m+t$, and $W\in\mathbb{R}^{|\mathcal{V}|\times H}$ is the output head mapping hidden states to vocabulary logits (Radford2018ImprovingLU; vaswani2017attention). Scaling inference of such models often requires distributing computation across multiple devices via tensor parallelism, which partitions per-layer parameters across devices and aggregates partial results with collectives such as AllReduce (HansenPalmus2024CommunicationCF; Li2024TPILLMS7). The per-token latency is then set by the critical path combining local compute and synchronizations.
Speculative decoding (SD). Speculative decoding augments a target model $f_{target}(\cdot\mid\cdot)$ with a computationally cheaper draft model $f_{draft}(\cdot\mid\cdot)$  (leviathan2023fast; chen2023accelerating). At step $t$, conditioned on the verified prefix $(x,y_{<t})$, the draft proposes a $\gamma$-token window
@@
Increasing draft capacity (larger $\gamma$, deeper/wider $f_{d}$) typically increases $\rho$ but also increases $T_{\text{draft}}$, while tiny drafts reduce $T_{\text{draft}}$ but suffer low $\rho$ (leviathan2023fast; chen2023accelerating).  Equation 5 exposes the core limitation: improvements in acceptance must compensate for the added draft latency, intrinsically coupling acceptance with latency.
@@ 3 Mirror Speculative Decoding
which is typically strongly correlated with the final distribution $p^{(N)}(\cdot\mid y_{<t},x)$  (Pal2023FutureLA). We designate an early-exit layer $\ell_{e}\in\{1,\dots,N-1\}$ and expose a low-bandwidth token channel:
@@ 3.2 Draft execution with Speculative Streaming
For the draft model $f_{d}$, we employ Speculative Streaming (SS) (bhendawade2024speculative), a speculative mechanism that verifies previously proposed tokens while generating new speculative tokens in the same forward pass using multi-stream attention. Applying SS to the target would modify its decoding dynamics and alter the final distribution $p^{(N)}(\cdot\mid y_{<t},x)$ (bhendawade2024speculative), breaking the lossless guarantee established in  Appendix B. In contrast, using SS on the draft accelerates speculation generation without changing acceptance semantics, since all commitments still require verification against the unchanged target. This design leverages SS precisely where it yields additional concurrency while preserving correctness (See  Appendix B).  Section D.2 illustrates the SS mechanism and compares draft-only speedups between vanilla and SS drafts.
@@
Within each Mirror-SD step, the draft must furnish a branch-complete speculative window of length $\gamma$ at the rendezvous ( Section 3.1). Under SS, a single draft internal step can emit $\eta_{j}\!\geq\!1$ tokens by verifying the prior proposal and predicting multiple future tokens in one pass (bhendawade2024speculative). Consequently, the number of draft steps $J$ required to materialize $\gamma$ tokens satisfies
@@ 3.3 Heterogeneous Sharding of Mirror-SD
The draft is trained with SPD architecture  (Kim2025SPDSD). We divide the $N_{\mathrm{d}}$ layers into two contiguous segments. Within each segment we instantiate $G_{D}$ parallel tracks; track $g\in\{1,\dots,G_{D}\}$ is pinned to NPU $g$ and advances through all layers of its segment using a resident weight shard. There is no inter-NPU traffic inside a segment (See  Figure 2). At the segment boundary, all tracks perform a single global synchronization to re-align tensor partitions, and a second synchronization occurs at the end of the forward pass to assemble full-width logits for the main and lookahead streams. Each internal draft step executes two all-reduce collectives on activation shards while weights remain sharded. This replaces per-layer synchronization with a fixed two-collective cost, reducing latency and enabling more parameters to be sharded across NPUs. In practice, this expands draft capacity and improves acceptance rates $\rho(\gamma;\phi,\theta)$ without increasing critical-path latency.
@@ 4.1 Evaluation protocol
Datasets and tasks. We integrate our approach with the open-source SpecBench framework (xia-etal-2024-unlocking) to ensure a fair, reproducible comparison against prior methods. SpecBench provides standardized prompts and pre/post-processing, sampling settings and released configs and seeds (xia-etal-2024-unlocking). We report results on multi-turn interactive conversation (MT Bench), translation, summarization, mathematical reasoning, machine translation and retrieval-augmented generation (RAG). Context and generation lengths follow the SpecBench protocol (xia-etal-2024-unlocking).
Models and baselines. We evaluate Mirror-SD on server-scale targets that are deployable in production inference stacks: Qwen3-14B and Qwen3-32B (qwen3technicalreport), Mistral-24B (mistral_small_3_2501_2025), and OPT-66B (zhang2022opt). For Qwen targets, we train a 0.6B-parameter draft with 2 segments and 8 tracks and deploy it on 8 NPUs as described in Section 3.3. For Mistral we train a 0.5B draft, and for OPT we train a 200M draft, both sharded as in Section 3.3 to optimize synchronization cost. All draft models are trained with SS objective described in (bhendawade2024speculative) on UltraChat  (ding2023enhancing). Across all target models, drafts are launched from the mid-layer early exit ($\frac{1}{2}$ of total depth) with top-$\kappa$=8 under batch size 1. Please refer to Appendix E for the effects of early-exit depth and $\kappa$. Baselines include vanilla SD, Medusa (Cai2024MedusaSL), Hydra (hydra), EAGLE 2/3 (eagle2; eagle3), Recycling (luo2024turning), PLD (saxena2023pld), SpS (gante2023assisted), REST (he-etal-2024-rest), and Lookahead (fu2023lookahead). All baselines have public implementations in SpecBench (xia-etal-2024-unlocking), and we use the corresponding implementations.
@@
Serving configuration and reproducibility. Target models are distributed across eight M2 Ultra GPUs using Megatron-style tensor parallelism (Section 3.3), while the draft runs on eight NPUs (apple_m2_ultra). All evaluations use a fixed batch size of 1 and speculative window length $\gamma{=}7$; please refer to Section D.1 for analysis of batching effects. The token channel transmits only the top-$\kappa$ token IDs and log-probabilities in bf16. For determinism, interconnects are pinned and frequency scaling is disabled. Timings include compute, collectives, and rendezvous overhead.
@@ 4.2 Tri-objective analysis with an MT-Bench diagnostic
Speculative decoding couples three quantities: the speculative window $\gamma$, the acceptance length $\mathbb{E}[A_{t}]=\gamma\,\rho(\gamma;\phi,\theta)$, and the per-speculation-step latency. In vanilla SD, enlarging $\gamma$ typically boosts acceptance but also increases draft construction time, yielding an upward-sloping latency curve. For Mirror-SD, the step latency follows the model in  Section 3.4 (Equation 10): as long as $T_{\text{draft}}^{\mathrm{gen}}(\gamma)\!\leq\!\Delta$ with $\Delta=T_{\text{target}}^{\,\ell_{e}+1:N}$, increasing $\gamma$ (and thus $\mathbb{E}[A_{t}]$) adds no marginal latency; once $T_{\text{draft}}^{\mathrm{gen}}(\gamma)\!>\!\Delta$, latency grows by the excess beyond $\Delta$. Acceptance semantics remain unchanged ( Appendix B). We validate these hypotheses on MT-Bench (bai-etal-2024-mt) by sweeping $\gamma$, measuring $\mathbb{E}[A_{t}]$ and the observed draft construction time, and comparing three methods that share the same target: (i) vanilla SD with autoregressive drafts from 12M to 1.7B parameters, (ii) Mirror-SD with a 0.60.6M draft, and (iii) Mirror-SD with a speculative-streaming draft (Section 3.2) of 0.60.6B.  Figure 3(a) places $\gamma$ on the x-axis, $\mathbb{E}[A_{t}]$ on the y-axis, and draft construction latency on the z-axis.
@@ Appendix A Related Works
The original speculative decoding paradigm accelerates autoregressive generation by pairing a small, fast draft model with a larger target model, which verifies proposed tokens (chen2023accelerating; leviathan2023fast). This approach achieves substantial wall-time savings whenever the draft is hardware-efficient and closely aligned with the target. Domain-specialized drafts trained via distillation further improve acceptance in task-specific settings (hong2025training). Recent variants explore parallelization strategies, such as batch-axis speculation (sun2023spectr) and tree-structured drafts (miao2023specinfer; spector2023accelerating), to raise acceptance rates and amortize draft cost.
@@
An alternative line of work removes the explicit draft model and equips the target itself with speculative capacity. Medusa predicts multiple tokens in parallel via extra heads (medusa), while Hydra enforces autoregressive coupling across those heads to raise acceptance (hydra). EAGLE introduces a dedicated speculation layer (eagle), with EAGLE-2 enabling dynamic tree retries (eagle2) and EAGLE-3 moving to token-level prediction with multi-layer fusion (eagle3). Prompt-lookup decoding (PLD) and Lookahead propose suffixes by retrieval rather than generation (saxena2023pld; fu2023lookahead), which is effective when prefix–continuation correlations are strong. Recycling reduces wasted work by reusing intermediate activations when speculative branches are invalidated, instead of recomputing full forwards (luo2024turning). Other recent advances include structured or retrieval-based decoding policies (yi2024generation; he-etal-2024-rest). Across the single-model designs, speculative capacity is integrated into the target stack, so larger or wider modules increase acceptance but still add work on the target’s critical path; by contrast, Mirror-SD runs draft and target on heterogeneous devices and overlaps draft within the target’s suffix window, converting added draft capacity into acceptance gains without inflating per-step latency proportionally.
@@
Beyond speculation, a range of methods accelerate inference by adapting compute during decoding. CALM (schuster2022confident) and related early-exit methods reduce cost by exiting tokens at shallow layers, while skip decoding (SkipDecode) mitigates key-value cache mismatch via position-dependent layer skipping. Mixture-of-Depths (MoD) (raposo2024mixture) routes only a subset of tokens through full blocks, yielding non-uniform FLOP allocation. Other strategies include token merging (bolya2023token) to reduce sequence length dynamically, adaptive span models (sukhbaatar2019adaptive) that learn context windows per token, and CoLT5 (ainslie2023colt5) which routes tokens through heavy or light pathways. More recently, M2R2 (bhendawade2025m2r2) introduces accelerated residual streams to improve early alignment and efficiency. Together, these approaches trade fixed per-token compute for dynamic allocation, complementing speculative decoding’s strategy of parallelizing token generation.